Jackie mixes some 35% acid solution with some 70% acid solution to make 100mL of a 42% acid solution. How much of the 35% acid solution did she use?
Answer by Maths68(1474)

Solution A
Amount = x ml
Concentration =35% =0.35
Solution B
Amount = 100-x ml
Concentration =70% = 0.7
Resultant Solution
Amount =100ml
Concentration =42%=0.42
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
(x)(0.35)+(0.7)(100-x)=100*.42
0.35x+70-0.7x=42
0.35x-0.7x=42-70
-0.35x=-28
-0.35x/-0.35=-28/-0.35
x=80

Jackie used 80ml of 35% acid solution and 20ml of 70% acid solution to make 100ml of a 42% acid solution.